ID |
Date |
Author |
Type |
Category |
Subject |
331
|
Fri Mar 10 16:44:13 2017 |
Gabriele | Optics | Daily Progress | Updated in vacuum periscope of CR0 |
This afternoon I removed the old periscope from CR0 and installed a new one with finely adjustable mount, like those in the new chamber. I realigned the optical lever to the horizontal refererence. |
330
|
Tue Mar 7 11:09:33 2017 |
Gabriele | General | Measurements | New MIT sample #3 |
2017-03-07
- 11:08am, in chamber CR0
- 11:09am, roughing pump on
- 11:23am turbo pump on
- 3:37pm, pumps stopped, the turbo is saturating CR1
2017-03-08
- 8:00am pumps restarted
- Excitations
-
Quiet time before excitation: 1173048953
Excitation broadband: 1173048986
Quiet time after excitation: 1173049018
-
Quiet time before excitation: 1173050248
Excitation broadband: 1173050280
Quiet time after excitation: 1173050312
-
Quiet time before excitation: 1173051542
Excitation broadband: 1173051574
Quiet time after excitation: 1173051606
-
Quiet time before excitation: 1173052836
Excitation broadband: 1173052868
Quiet time after excitation: 1173052900
-
Quiet time before excitation: 1173054130
Excitation broadband: 1173054163
Quiet time after excitation: 1173054195
-
Quiet time before excitation: 1173055425
Excitation broadband: 1173055457
Quiet time after excitation: 1173055489
2017-03-09
|
329
|
Tue Mar 7 11:01:58 2017 |
Gabriele | General | Measurements | S1600469 S1600471 S1600481 S1600483 |
2017-03-07
- 10:57am, in chamber
- S1600469 in CR1
- S1600471 in CR2
- S1600481 in CR3
- S1600483 in CR4
- 11:02am, roughing pump on
- 11:11am, turbo pump on
- Excitations
-
Quiet time before excitation: 1172968827
Excitation broadband: 1172968862
Quiet time after excitation: 1172968927
-
Quiet time before excitation: 1172970757
Excitation broadband: 1172970792
Quiet time after excitation: 1172970857
-
Quiet time before excitation: 1172972687
Excitation broadband: 1172972723
Quiet time after excitation: 1172972788
-
Quiet time before excitation: 1172974618
Excitation broadband: 1172974653
Quiet time after excitation: 1172974718
-
Quiet time before excitation: 1172976548
Excitation broadband: 1172976583
Quiet time after excitation: 1172976648
-
Quiet time before excitation: 1172978478
Excitation broadband: 1172978513
Quiet time after excitation: 1172978578
-
Quiet time before excitation: 1172980408
Excitation broadband: 1172980443
Quiet time after excitation: 1172980508
-
Quiet time before excitation: 1172982338
Excitation broadband: 1172982373
Quiet time after excitation: 1172982438
2017-03-08
- 8:00am pumps off, gate valve closed
|
328
|
Fri Mar 3 13:01:59 2017 |
Gabriele | General | Measurements | New MIT sample #2 |
2017-03-03
- 1:00pm, in chamber, realigned optical lever to get as close as possible to the edge of the disk
- 1:02pm, roughing pump on
- 1:14pm, turbo pump on
- Excite
-
Quiet time before excitation: 1172633160
Excitation broadband: 1172633192
Quiet time after excitation: 1172633224
-
Quiet time before excitation: 1172634454
Excitation broadband: 1172634486
Quiet time after excitation: 1172634518
-
Quiet time before excitation: 1172635748
Excitation broadband: 1172635781
Quiet time after excitation: 1172635813
-
Quiet time before excitation: 1172637043
Excitation broadband: 1172637075
Quiet time after excitation: 1172637107
-
Quiet time before excitation: 1172638337
Excitation broadband: 1172638369
Quiet time after excitation: 1172638401
-
Quiet time before excitation: 1172639631
Excitation broadband: 1172639663
Quiet time after excitation: 1172639695
2017-03-06
|
327
|
Fri Mar 3 13:00:50 2017 |
Gabriele | General | Measurements | S1600453 S1600490 |
2017-03-03
- 12:45pm, in chamber
- S1600453 in CR3
- S1600490 in CR4
- 12:45pm roughing pump on
- 12:56pm, turbo pump on
- Excitation
-
Quiet time before excitation: 1172620986
Excitation broadband: 1172621021
Quiet time after excitation: 1172621086
-
Quiet time before excitation: 1172622916
Excitation broadband: 1172622951
Quiet time after excitation: 1172623016
-
Quiet time before excitation: 1172624846
Excitation broadband: 1172624881
Quiet time after excitation: 1172624946
-
Quiet time before excitation: 1172626776
Excitation broadband: 1172626811
Quiet time after excitation: 1172626876
-
Quiet time before excitation: 1172628706
Excitation broadband: 1172628742
Quiet time after excitation: 1172628807
-
Quiet time before excitation: 1172630638
Excitation broadband: 1172630673
Quiet time after excitation: 1172630738
-
Quiet time before excitation: 1172632568
Excitation broadband: 1172632603
Quiet time after excitation: 1172632668
-
Quiet time before excitation: 1172634498
Excitation broadband: 1172634534
Quiet time after excitation: 1172634599
2017-03-06
- 10:19am, gate valve closed, pumps off
|
326
|
Thu Mar 2 11:27:47 2017 |
Gabriele | General | Noise hunting | Turbo pump of CR0 pollutes CR1 |
The turbo pump of the CR0 chamber runs at 833 Hz. It causes vibrations that pollute the measurements in the CR1-4 chamber. In particular CR1 is extremely sensitive and the line is highly up-converted. It's not clean why CR1 is more sensitive than CR2-3-4.

|
325
|
Thu Mar 2 09:45:58 2017 |
Gabriele | General | Measurements | S1600456 S1600458 S1600461 S1600491 |
2017-03-02
- 9:31am, in chamber, removed ND from path 2-4
- S1600456 in CR1
- S1600458 in CR2
- S1600461 in CR3
- S1600491 in CR4
- 9:45am roughing pump on
- 9:55am turbo pump on
- Excitations
-
Quiet time before excitation: 1172530103
Excitation broadband: 1172530139
Quiet time after excitation: 1172530204
-
Quiet time before excitation: 1172532034
Excitation broadband: 1172532069
Quiet time after excitation: 1172532134
-
Quiet time before excitation: 1172533964
Excitation broadband: 1172533999
Quiet time after excitation: 1172534064
-
Quiet time before excitation: 1172535894
Excitation broadband: 1172535929
Quiet time after excitation: 1172535994
-
Quiet time before excitation: 1172537824
Excitation broadband: 1172537859
Quiet time after excitation: 1172537924
-
Quiet time before excitation: 1172539754
Excitation broadband: 1172539789
Quiet time after excitation: 1172539854
-
Quiet time before excitation: 1172541684
Excitation broadband: 1172541719
Quiet time after excitation: 1172541784
-
Quiet time before excitation: 1172543614
Excitation broadband: 1172543649
Quiet time after excitation: 1172543714
2017-03-03
- 9:30am, gate valve closed, pumps stopped, venting
|
324
|
Wed Mar 1 10:10:47 2017 |
Gabriele | General | Measurements | S1600469 S1600471 S1600481 S1600483 |
2017-03-01
- 10:10am, in chamber, added neutral densities
- S1600469 in CR1
- S1600471 in CR2
- S1600481 in CR3
- S1600483 in CR4
- 10:12am, roughing pump on, clamps on
- 10:22am, turbo pump on
- Excitations:
-
Quiet time before excitation: 1172438462
Excitation broadband: 1172438497
Quiet time after excitation: 1172438562
-
Quiet time before excitation: 1172440392
Excitation broadband: 1172440427
Quiet time after excitation: 1172440492
-
Quiet time before excitation: 1172442322
Excitation broadband: 1172442357
Quiet time after excitation: 1172442422
-
Quiet time before excitation: 1172444252
Excitation broadband: 1172444287
Quiet time after excitation: 1172444352
-
Quiet time before excitation: 1172446182
Excitation broadband: 1172446217
Quiet time after excitation: 1172446282
-
Quiet time before excitation: 1172448112
Excitation broadband: 1172448148
Quiet time after excitation: 1172448213
-
Quiet time before excitation: 1172450044
Excitation broadband: 1172450079
Quiet time after excitation: 1172450144
-
Quiet time before excitation: 1172451974
Excitation broadband: 1172452009
Quiet time after excitation: 1172452074
2017-03-02
- 9:00am, vented, samples removed
|
323
|
Tue Feb 28 14:20:50 2017 |
Gabriele | General | Measurements | New MIT sample #1 |
2017-02-28
- 2:19pm, installed in CR0, balanced, with 1/2" inserts facing down
- It looks dusty and there is a fingerprint on the edge. I didn't clean it
- 2:21pm, roughing pump on
- 2:35pm, turbo pump on
- Excitations:
-
Quiet time before excitation: 1172377933
Excitation broadband: 1172377965
Quiet time after excitation: 1172378027
-
Quiet time before excitation: 1172381657
Excitation broadband: 1172381689
Quiet time after excitation: 1172381751
-
Quiet time before excitation: 1172385381
Excitation broadband: 1172385414
Quiet time after excitation: 1172385476
-
Quiet time before excitation: 1172389106
Excitation broadband: 1172389138
Quiet time after excitation: 1172389200
-
Quiet time before excitation: 1172392830
Excitation broadband: 1172392863
Quiet time after excitation: 1172392925
-
Quiet time before excitation: 1172396555
Excitation broadband: 1172396587
Quiet time after excitation: 1172396649
2017-03-01
2017-03-02
- 9:00am. The laser beam was probably hitting one of the glued inserts. Checked that te disk is not touching, moved the ESD a bit lower
- 9:04am, roughing pump on
- 9:17am, turbo pump on
- 1:50pm, pumps off
|
322
|
Tue Feb 28 11:22:35 2017 |
Gabriele | General | Measurements | S1600463 S1600466 S1600475 S1600476 |
2017-02-28
- 11:20am, in chamber
- S1600463 in CR1
- S1600466 in CR2
- S1600475 in CR3
- S1600476 in CR4
- 2:10pm, roughing pump on
- 2:20pm, turbo pump on
- Excitations
-
Quiet time before excitation: 1172366632
Excitation broadband: 1172366667
Quiet time after excitation: 1172366732
-
Quiet time before excitation: 1172373962
Excitation broadband: 1172373997
Quiet time after excitation: 1172374062
-
Quiet time before excitation: 1172381292
Excitation broadband: 1172381327
Quiet time after excitation: 1172381392
-
Quiet time before excitation: 1172388622
Excitation broadband: 1172388657
Quiet time after excitation: 1172388722
2017-03-01
|
321
|
Mon Feb 27 13:13:05 2017 |
Gabriele | General | Measurements | S1600490 |
2017-02-27
- 12:59pm, installed in chamber CR0, balanced
- 1:00pm, roughing pump on
- 1:13pm, turbo pump on
- 3:04pm, stopped pumps
|
320
|
Mon Feb 27 08:42:05 2017 |
Gabriele | General | Measurements | S1600453 S1600490 |
2017-02-27
- 8:40am, in chamber, balanced
- S1600453 in CR1
- S1600491 in CR2 (i meant to install 490, but got confused)
- 10:23am, roughing pump on, clamps on
- 10:32am, turbo pump on
- 3:01pm, something is wrong with CR1 (the 666 Hz line due to the turbo pump is very large and saturates the QPD signals). Venting
- 4:11pm in chamber
- S1600453 in CR1
- S1600490 in CR2
- 4:12pm roughing pump on, clamps on
- 4:20pm turbo pump on
- Excitations
-
Quiet time before excitation: 1172290836
Excitation broadband: 1172290871
Quiet time after excitation: 1172290936
-
Quiet time before excitation: 1172292766
Excitation broadband: 1172292801
Quiet time after excitation: 1172292866
-
Quiet time before excitation: 1172294696
Excitation broadband: 1172294731
Quiet time after excitation: 1172294796
-
Quiet time before excitation: 1172296626
Excitation broadband: 1172296661
Quiet time after excitation: 1172296726
-
Quiet time before excitation: 1172298556
Excitation broadband: 1172298591
Quiet time after excitation: 1172298656
-
Quiet time before excitation: 1172300486
Excitation broadband: 1172300522
Quiet time after excitation: 1172300587
-
Quiet time before excitation: 1172302417
Excitation broadband: 1172302452
Quiet time after excitation: 1172302518
-
Quiet time before excitation: 1172304348
Excitation broadband: 1172304383
Quiet time after excitation: 1172304448
2017-02-28
|
319
|
Fri Feb 24 14:44:22 2017 |
Gabriele | General | General | Disk model |
It's possible to build an analytical model of the resonant frequencies of a simple thin disk. For example, see J. Sound and Vibrations 188, 685 (1995), section 2
The solutions are given in term of Bessel functions:
![W_{mn}(r, \theta)=\left[ J_m(\lambda_{mn} r/a) + C_{mn} I_m(\lambda_{mn} r/a)\right] \cos m \theta](https://latex.codecogs.com/gif.latex?W_%7Bmn%7D%28r%2C%20%5Ctheta%29%3D%5Cleft%5B%20J_m%28%5Clambda_%7Bmn%7D%20r/a%29%20+%20C_%7Bmn%7D%20I_m%28%5Clambda_%7Bmn%7D%20r/a%29%5Cright%5D%20%5Ccos%20m%20%5Ctheta)
where J is a Bessel function of the first kind, and I a modified Bessel function of the first kind, a is the disk radius.
The coefficient Cmn and the eigenvalue can be found as solution of the following two equations

Then the eigenfrequencies are given by

where rho is the material density, h the disk thickness and D the flexural rigidity

where E is the material Young's modulus and nu the Poisson's ratio.
From all those results we can conclude that the frequency scaling with respect to disk radius and thickness are very simple:
- the frequencies scale linearly with the thickness
- the frequencies scale inversely proportionally to the square of the radius
Also, the frequencies scales as sqrt(E/rho)

The dependency on the Poisson's ratio is more complex since nu is involved in the eigenfrequency equation shown above.
Unfortunately the thin disk model does not exactly match the COMSOL results: deviations of few tens of Hz are present, probaly due to the thin disk approximation. The COMSOL model is more accurate to match the experimental frequencies.
However, I checked that the eigenfrequencies predicted by COMSOL also scales as predicted with thickness and radius.
Using the measurements on the six samplex we got from Mark Optics, after annealing, I was able to tune the COMSOL model to fit all measured frequencies within 6 Hz. I chose to change the disk thickness (since diameter and Young'r modulus are degenerate) and the Poisson's ratio.
Here is an example of the difference between the measured and modeled frequencies:

The table below summarizes the best fit for each of the disks
SN |
Thickness [mm] |
Poisson's ratio |
S1600519 |
1.0194 |
0.1669 |
S1600520 |
1.0186 |
0.1663 |
S1600521 |
1.0217 |
0.1669 |
S1600522 |
1.0223 |
0.1635 |
S1600523 |
1.0227 |
0.1655 |
S1600524 |
1.0209 |
0.1654 |
Since the material is the same, I would expect the Poisson's ratio to be constant. So for future modeling I'm using the average of the values above: 0.166 |
318
|
Fri Feb 24 13:21:03 2017 |
Gabriele | General | Measurements | Q of substrates with polished edges |
Below the measured Q values of the six Suprasil 313 substrates, with polished edges, after annealing at 900 degrees C for 9 hours.
 
 
 
They are all consistently at a level of 100e6 at 1 KHz and showing a similar frequency dependency going down to a about 30e6 at 20-30 kHz. The only exception is S1600522 which has clearly lower Q values at all frequencies.
Below a comparison of the substrate loss angles compared wiht Steve Penn's model:

|
317
|
Fri Feb 24 11:05:58 2017 |
Gabriele | General | Measurements | S1600456 S1600458 S1600461 S1600491 |
2017-02-24
- 10:50am, in chamber, balanced, added neutral density to laser on the 1/3 path
- S1600456 in CR1
- S1600458 in CR2
- S1600461 in CR3
- S1600491 in CR4
- 10:56am, roughing pump on, clamps on
- 11:05am, turbo pump on
- Excitations
-
Quiet time before excitation: 1172009197
Excitation broadband: 1172009232
Quiet time after excitation: 1172009297
-
Quiet time before excitation: 1172011127
Excitation broadband: 1172011162
Quiet time after excitation: 1172011227
-
Quiet time before excitation: 1172013057
Excitation broadband: 1172013092
Quiet time after excitation: 1172013157
-
Quiet time before excitation: 1172014987
Excitation broadband: 1172015022
Quiet time after excitation: 1172015087
-
Quiet time before excitation: 1172016917
Excitation broadband: 1172016952
Quiet time after excitation: 1172017017
-
Quiet time before excitation: 1172018847
Excitation broadband: 1172018882
Quiet time after excitation: 1172018947
-
Quiet time before excitation: 1172020777
Excitation broadband: 1172020812
Quiet time after excitation: 1172020877
-
Quiet time before excitation: 1172022707
Excitation broadband: 1172022742
Quiet time after excitation: 1172022807
2017-02-27
|
316
|
Thu Feb 23 10:54:41 2017 |
Gabriele | General | Measurements | S1600523 S1600524 |
2017-02-23
- 10:35am in chamber, balanced
- S1600523 in bay 1
- S1600524 in bay 2
- 10:45am, roughing pump on, clamps on
- 10:55am, turbo pump on
- Excitations:
-
Quiet time before excitation: 1171925035
Excitation broadband: 1171925070
Quiet time after excitation: 1171925085
-
Quiet time before excitation: 1171935915
Excitation broadband: 1171935950
Quiet time after excitation: 1171935965
-
Quiet time before excitation: 1171946795
Excitation broadband: 1171946830
Quiet time after excitation: 1171946845
-
Quiet time before excitation: 1171957675
Excitation broadband: 1171957710
Quiet time after excitation: 1171957725
2017-02-24
- 8:40am, gate valve close, pumps off, venting
|
315
|
Wed Feb 22 15:11:59 2017 |
Gabriele | General | Measurements | S1600519 S1600520 S1600521 S1600522 |
2017-02-22
- 2:59pm, in chamber, balanced, removed all neutral densities
- 3:02pm, roughing pump on, clamps on
- 3:11pm, turbo pump on
- Excitations
-
Quiet time before excitation: 1171847838
Excitation broadband: 1171847873
Quiet time after excitation: 1171847888
-
Quiet time before excitation: 1171858718
Excitation broadband: 1171858754
Quiet time after excitation: 1171858769
-
Quiet time before excitation: 1171869599
Excitation broadband: 1171869634
Quiet time after excitation: 1171869649
-
Quiet time before excitation: 1171880479
Excitation broadband: 1171880514
Quiet time after excitation: 1171880530
2017-02-23
- 10:22am, gate valve closed, pumps off, venting
|
314
|
Wed Feb 22 13:56:15 2017 |
Gabriele | Electronics | Configuration | Software improvements |
- created a single interface to move the picomotors in front of QPD1 ... QPD4
- the command
qpdcenter will open an interactive python shell to control the picomotors
- two scripts for automatic excitations are saved in
/opt/rtcds/userapps/CyMAC/src/auto_excite*.py
- they must run on the cymac. But they can be started from the control station using the commands
autoexcite0 and autoexcite14
- before starting the commands, set the parameters in the file. The two commands
edit_autoexcite0 and edit_autoexcite14 will open the correct files
- a log file of the excitation times will be saved
- the scripts will check if there is an excitation going
- the two commands
noise0 and noise14 can be used to start awggui to inject the correct noise
- updated the autocentering scripts, so that only one instance can run at a time.
- launch them with the commands
autocenter0 and autocenter14
|
313
|
Sun Feb 19 11:46:26 2017 |
Gabriele | General | Measurements | S1600458 S1600474 S1600491 |
2017-02-19
- 11:28am an chamber:
- S1600458 in CR1
- S1600474 in CR2
- S1600491 in CR3
- empty CR4
- 11:46am roughing pump on, clamps on
- 11:57am turbo pump on
- Excitations:
-
Quiet time before excitation: 1171579753
Excitation (broad band) at 1171579788 (60 s)
Quiet time after excitation: 1171579854
-
Quiet time before excitation: 1171581684
Excitation (broad band) at 1171581719 (60 s)
Quiet time after excitation: 1171581784
-
Quiet time before excitation: 1171583614
Excitation (broad band) at 1171583649 (60 s)
Quiet time after excitation: 1171583714
-
Quiet time before excitation: 1171585545
Excitation (broad band) at 1171585580 (60 s)
Quiet time after excitation: 1171585645
-
Quiet time before excitation: 1171587475
Excitation (broad band) at 1171587510 (60 s)
Quiet time after excitation: 1171587575
-
Quiet time before excitation: 1171589405
Excitation (broad band) at 1171589440 (60 s)
Quiet time after excitation: 1171589505
-
Quiet time before excitation: 1171591335
Excitation (broad band) at 1171591370 (60 s)
Quiet time after excitation: 1171591435
-
Quiet time before excitation: 1171593265
Excitation (broad band) at 1171593300 (60 s)
Quiet time after excitation: 1171593365
2017-02-21
- 10:30am, gate valve closed, pumps stopped, venting
|
312
|
Sat Feb 18 15:04:00 2017 |
Gabriele | General | Measurements | S1600481 S1600483 S1600469 S1600471 |
2017-02-18
- 9:54am in chamber, balanced
- S1600469 in CR1
- S1600471 in CR2
- S1600481 in CR3
- S1600483 in CR4
- 10:00am, roughing pump on, clamps on
- 10:13am, turbo pump on
- Excitations
-
Quiet time before excitation: 1171492847
Excitation (broad band) at 1171492882 (60 s)
Quiet time after excitation: 1171492947
-
Quiet time before excitation: 1171494777
Excitation (broad band) at 1171494813 (60 s)
Quiet time after excitation: 1171494878
-
Quiet time before excitation: 1171496708
Excitation (broad band) at 1171496743 (60 s)
Quiet time after excitation: 1171496808
-
Quiet time before excitation: 1171498638
Excitation (broad band) at 1171498673 (60 s)
Quiet time after excitation: 1171498738
-
Quiet time before excitation: 1171500568
Excitation (broad band) at 1171500603 (60 s)
Quiet time after excitation: 1171500668
-
Quiet time before excitation: 1171502498
Excitation (broad band) at 1171502533 (60 s)
Quiet time after excitation: 1171502598
-
Quiet time before excitation: 1171504428
Excitation (broad band) at 1171504463 (60 s)
Quiet time after excitation: 1171504528
-
Quiet time before excitation: 1171506358
Excitation (broad band) at 1171506393 (60 s)
Quiet time after excitation: 1171506458
-
More excitations:
-
Quiet time before excitation: 1171509354
Excitation (broad band) at 1171509389 (60 s)
Quiet time after excitation: 1171509454
-
Quiet time before excitation: 1171511284
Excitation (broad band) at 1171511319 (60 s)
Quiet time after excitation: 1171511384
-
Quiet time before excitation: 1171513214
Excitation (broad band) at 1171513249 (60 s)
Quiet time after excitation: 1171513314
-
Quiet time before excitation: 1171515144
Excitation (broad band) at 1171515179 (60 s)
Quiet time after excitation: 1171515244
-
Quiet time before excitation: 1171517074
Excitation (broad band) at 1171517109 (60 s)
Quiet time after excitation: 1171517174
-
Quiet time before excitation: 1171519004
Excitation (broad band) at 1171519039 (60 s)
Quiet time after excitation: 1171519104
-
Quiet time before excitation: 1171520934
Excitation (broad band) at 1171520969 (60 s)
Quiet time after excitation: 1171521034
-
Quiet time before excitation: 1171522864
Excitation (broad band) at 1171522899 (60 s)
Quiet time after excitation: 1171522964
2017-02-19
- 11:11am, gate valve closed, pumps stopped, venting
|
311
|
Fri Feb 17 13:44:47 2017 |
Gabriele | Optics | Daily Progress | Swapped picomotor and re-alignment |
Since I had recurrent problems with the picomotors used for QPD3, I swapped them with another Newport motorized mirror that was previously used in the Crackle1 experiment. This is the same model used for the other three QPD centering. Everything looks to be working fine now.
I also realigned all optical levers and swapped out an iris with a smaller one, to avoid beam clipping. All beam paths look clear now. |
310
|
Thu Feb 16 10:19:33 2017 |
Gabriele | General | Measurements | S1600453 S1600456 S1600461 S1600490 |
2017-02-16
- 10:17am in chamber
- S1600453 in CR1
- S1600456 in CR2
- S1600461 in CR3
- S1600490 in CR4
- Notes:
- added ND03 on path to CR1 and CR3, since QPD were saturating
- there is something wrong with CR1, the spectrum is very noisy. I checked that there is no clipping of the beam. I also tried to move the spot on the disk, without change. Maybe the disk is touching?
- 10:18am, roughing pump on, clamps on
- 10:21am, turbo pump on
- Excitations: (NOTE: bay 1 noise spectrum is very bad. The autocetering of bay 3 failed, so the QPD was not centered for most of the time)
-
Quiet time before excitation: 1171326574
Excitation (broad band) at 1171326609 (60 s)
Quiet time after excitation: 1171326674
-
Quiet time before excitation: 1171328504
Excitation (broad band) at 1171328539 (60 s)
Quiet time after excitation: 1171328604
-
Quiet time before excitation: 1171330434
Excitation (broad band) at 1171330469 (60 s)
Quiet time after excitation: 1171330534
-
Quiet time before excitation: 1171332364
Excitation (broad band) at 1171332399 (60 s)
Quiet time after excitation: 1171332465
-
Quiet time before excitation: 1171334295
Excitation (broad band) at 1171334330 (60 s)
Quiet time after excitation: 1171334395
-
Quiet time before excitation: 1171336225
Excitation (broad band) at 1171336260 (60 s)
Quiet time after excitation: 1171336325
-
Quiet time before excitation: 1171338155
Excitation (broad band) at 1171338191 (60 s)
Quiet time after excitation: 1171338256
-
Quiet time before excitation: 1171340086
Excitation (broad band) at 1171340121 (60 s)
Quiet time after excitation: 1171340186
2017-02-17
- 8:43am gate valve closed, pump stopped, venting slowly
- the measurements carried out last night didn't look very good, mostly because it seems that the picomotor of QPD3 got crazy and was shaking too much the bench
- 12:19pm, balanced again, roughing pump on. Added ND03 and ND04 to lasers to avoid QPD saturations
- 12:49pm, turbo pump on
- Excitations:
-
Quiet time before excitation: 1171425425
Excitation (broad band) at 1171425460 (60 s)
Quiet time after excitation: 1171425525
-
Quiet time before excitation: 1171427355
Excitation (broad band) at 1171427390 (60 s)
Quiet time after excitation: 1171427455
-
Quiet time before excitation: 1171429285
Excitation (broad band) at 1171429320 (60 s)
Quiet time after excitation: 1171429385
-
Quiet time before excitation: 1171431215
Excitation (broad band) at 1171431250 (60 s)
Quiet time after excitation: 1171431315
-
Quiet time before excitation: 1171433145
Excitation (broad band) at 1171433180 (60 s)
Quiet time after excitation: 1171433245
-
Quiet time before excitation: 1171435075
Excitation (broad band) at 1171435110 (60 s)
Quiet time after excitation: 1171435176
-
Quiet time before excitation: 1171437006
Excitation (broad band) at 1171437041 (60 s)
Quiet time after excitation: 1171437106
-
Quiet time before excitation: 1171438936
Excitation (broad band) at 1171438971 (60 s)
Quiet time after excitation: 1171439036
2017-02-18
- 9:34am, gate valve closed, pumps stopped, venting
|
309
|
Thu Feb 16 10:07:59 2017 |
Gabriele | General | Measurements | S1600491 |
2017-02-16
- 10:06am, in chamber CR0, balanced
- 10:07am, roughing pump on
- 10:31am, turbo pump on
- Excitations
-
Quiet time before excitation: 1171326606
Excitation (broad band) at 1171326638 (60 s)
Quiet time after excitation: 1171326700
-
Quiet time before excitation: 1171328530
Excitation (broad band) at 1171328562 (60 s)
Quiet time after excitation: 1171328624
-
Quiet time before excitation: 1171330454
Excitation (broad band) at 1171330487 (60 s)
Quiet time after excitation: 1171330549
-
Quiet time before excitation: 1171332379
Excitation (broad band) at 1171332411 (60 s)
Quiet time after excitation: 1171332473
-
Quiet time before excitation: 1171334303
Excitation (broad band) at 1171334335 (60 s)
Quiet time after excitation: 1171334397
-
Quiet time before excitation: 1171336227
Excitation (broad band) at 1171336259 (60 s)
Quiet time after excitation: 1171336321
-
Quiet time before excitation: 1171338151
Excitation (broad band) at 1171338184 (60 s)
Quiet time after excitation: 1171338246
-
Quiet time before excitation: 1171340076
Excitation (broad band) at 1171340108 (60 s)
Quiet time after excitation: 1171340170
2017-02-17
|
308
|
Thu Feb 16 09:47:29 2017 |
Gabriele | General | Measurements | Summary of coating runs carried out at Montreal |
Substrate temperature |
Serial numbers |
room temperature |
457 458 490 492 |
150 C |
489 491 |
250 C |
447 456 |
400 C |
459 461 |
480 C |
448 453 |
|
307
|
Tue Feb 14 15:50:23 2017 |
Gabriele | General | Measurements | S1600519 S1600520 S1600521 S1600522 |
New substrates from Mark Optics, round, with polished edges
2017-02-14
2017-02-15
- 3:16pm, gate valve closed, pump stopped, venting
|
306
|
Tue Feb 14 14:47:00 2017 |
Gabriele | General | Measurements | S1600495 |
2017-02-14
- 2:32pm, in chamber (CR0), balanced
- 2:33pm roughing pump on
- 2:45pm, turbo pump on
- Excitations:
-
Quiet time before excitation: 1171161389
Excitation (broad band) at 1171161421 (60 s)
Quiet time after excitation: 1171161483
-
Quiet time before excitation: 1171172313
Excitation (broad band) at 1171172345 (60 s)
Quiet time after excitation: 1171172407
-
Quiet time before excitation: 1171183237
Excitation (broad band) at 1171183269 (60 s)
Quiet time after excitation: 1171183331
-
Quiet time before excitation: 1171194161
Excitation (broad band) at 1171194193 (60 s)
Quiet time after excitation: 1171194255
-
Quiet time before excitation: 1171205085
Excitation (broad band) at 1171205117 (60 s)
Quiet time after excitation: 1171205179
2017-02-15
- 3:19pm, pumps stopped, venting
|
305
|
Tue Feb 14 10:55:13 2017 |
Gabriele | General | Daily Progress | Something fishy with bay 4 of the new setup |
I did two set of measurements with the new coated samples from Montreal. I reshuffled the position in the two measurements. In both cases, the measurement being performed in bay 4 was bad, in the sense that it was very hard to see excited modes. Since the two measurements were carried out with two different disks, it's clear it is a problem with that setup.
SOLVED: there was a connection problem for the DAC output signal controlling the switch
REALLY SOLVED: it was not a cabling issue. The power supply for the switching box had the current limiter on: when all four switches are closed, the box drain about 270mA, which is more than the limit of 250mA. Therefore the power supply voltage dropped and only three switches were actually closed. I switched the power supply to 500mA range and maxed the current limit. Now all four switches are working properly |
304
|
Sat Feb 11 16:22:26 2017 |
Gabriele | Mechanics | Daily Progress | Translation stage to move retaining rings |
The ring motion up and down was not very smooth, again due to friction on the centering pins.
So, after centering the rings using the pins and securing the rings to the translation stage, I removed all pins.
Now the motion up and down is very smooth.
I still have to fine tune the amount of steps that are needed to go up and down.
However, initial tests don't show a good repeatability of the positioning. My main suspect is that the vibration caused by the picomotor cause the disks to slip on the silicon lens. Indeed, when the disks are sitting on the rings, one can clearly hear them "rattle".
Quote: |
This afternoon I installed the picomotor and the translation stage that will be used to move the retaining rings up and down. No partciular problem: I only had to add some small aluminum foil shims between the ear of some rings and the square plate, to make the rings as horizontal as possible.
I tested the motion: with 300000 steps it's possible to move the rings all the way from the parked (down) position, to the up position. I also checked that when the rings are up, I can place four substarates and they fall properly into the alignment groove. Since the maximum speed of the picomotor is 2000 steps/s, it takes 150 seconds to move up and down the ring.
Finally, positive steps means that the rings are moving up, negative that they're moving down.
I raeligned the optical levers to the position I obtained by centering the samples with the rings. I haven't tested the repeatability yet.
|
|
303
|
Fri Feb 10 17:16:28 2017 |
Gabriele | Mechanics | Daily Progress | Translation stage to move retaining rings |
This afternoon I installed the picomotor and the translation stage that will be used to move the retaining rings up and down. No partciular problem: I only had to add some small aluminum foil shims between the ear of some rings and the square plate, to make the rings as horizontal as possible.
I tested the motion: with 300000 steps it's possible to move the rings all the way from the parked (down) position, to the up position. I also checked that when the rings are up, I can place four substarates and they fall properly into the alignment groove. Since the maximum speed of the picomotor is 2000 steps/s, it takes 150 seconds to move up and down the ring.
Finally, positive steps means that the rings are moving up, negative that they're moving down.
I raeligned the optical levers to the position I obtained by centering the samples with the rings. I haven't tested the repeatability yet.
 
|
302
|
Thu Feb 9 14:49:51 2017 |
Gabriele | General | Measurements | S1600489 S1600490 S1600491 S1600492 |
2017-02-09
- 2:30pm, samples installed in the chamber:
- S1600489 in bay 1
- S1600490 in bay 2
- S1600491 in bay 3
- S1600492 in bay 4
- 2:35pm, roughing pump on
- 2:45pm, pressure is <1 Torr, turbo pump on
- 4:19pm, pressure is 7.2e-6:
- but I forgot to clamp the lid!
- added four clamps at 4:22pm
- pressure seems to be going down again
- Excitations
-
Quiet time before excitation: 1170737383
Excitation (broad band) at 1170737418 (60 s)
Quiet time after excitation: 1170737483
-
Quiet time before excitation: 1170739313
Excitation (broad band) at 1170739349 (60 s)
Quiet time after excitation: 1170739414
-
Quiet time before excitation: 1170741244
Excitation (broad band) at 1170741279 (60 s)
Quiet time after excitation: 1170741344
-
Quiet time before excitation: 1170743174
Excitation (broad band) at 1170743210 (60 s)
Quiet time after excitation: 1170743275
-
Quiet time before excitation: 1170745105
Excitation (broad band) at 1170745140 (60 s)
Quiet time after excitation: 1170745205
-
Quiet time before excitation: 1170747035
Excitation (broad band) at 1170747070 (60 s)
Quiet time after excitation: 1170747135
-
Quiet time before excitation: 1170748965
Excitation (broad band) at 1170749000 (60 s)
Quiet time after excitation: 1170749065
-
Quiet time before excitation: 1170750896
Excitation (broad band) at 1170750931 (60 s)
Quiet time after excitation: 1170750996
2017-02-10
- 11:37am, gate valve closed, pumps off, venting
2017-10-13
- 8:30am, in chamber, balanced
- S1600492 in bay 1
- S1600491 in bay 2
- S1600490 in bay 3
- S1600489 in bay 4
- 8:34am roughing pump on, clamps on lid
- 8:46am turbo pump on
- Excitations
-
Quiet time before excitation: 1171048567
Excitation (broad band) at 1171048602 (60 s)
Quiet time after excitation: 1171048667
-
Quiet time before excitation: 1171049897
Excitation (broad band) at 1171049933 (60 s)
Quiet time after excitation: 1171049997
-
Quiet time before excitation: 1171051227
Excitation (broad band) at 1171051262 (60 s)
Quiet time after excitation: 1171051327
-
Quiet time before excitation: 1171052557
Excitation (broad band) at 1171052592 (60 s)
Quiet time after excitation: 1171052657
-
Quiet time before excitation: 1171053887
Excitation (broad band) at 1171053922 (60 s)
Quiet time after excitation: 1171053987
-
Quiet time before excitation: 1171055217
Excitation (broad band) at 1171055252 (60 s)
Quiet time after excitation: 1171055317
-
Quiet time before excitation: 1171056547
Excitation (broad band) at 1171056582 (60 s)
Quiet time after excitation: 1171056647
-
Quiet time before excitation: 1171057877
Excitation (broad band) at 1171057912 (60 s)
Quiet time after excitation: 1171057977
- Chamber opened at about 2:30pm, samples out
|
301
|
Thu Feb 9 11:35:27 2017 |
Gabriele | Optics | Configuration | QPD calibration |
This morning I measured the beam profiles at the QPD positions. As expected, the beams are not gaussian there, since we are seeing the interference of the reflections from the two faces of the disks. Nevertheless, the measured sizes are
|
Beam radius [microns] |
QPD1 |
302 |
QPD2 |
338 |
QPD3 |
270 |
QPD4 |
401 |
Here are the images of the beams:
 
 
The arm lever lengths (distance from the disk surface to the QPD) can be estimated from the optical drawing:
|
Distance disk to QPD [mm] |
QPD1 |
1540 |
QPD2 |
1813 |
QPD3 |
1587 |
QPD4 |
1652 |
The optical gain (normalized QPD signal over disk surface angle) is given by

|
Gain 1/rad |
QPD1 |
16300 |
QPD2 |
17100 |
QPD3 |
18800 |
QPD4 |
13150 |
The uncertaint in those numbers is quite large, both from the beam size and from the arm lever. Anyhow, I'm using the average to have a rough calibration of the QPD signals in terms of disk surface angular motion. The value I plugged in into the X_NORM filter banks is 16300 /rad. Therefore now the signals *NORM_OUT_DQ are calibrated in radians (deflection of the disk surface).

Finally, I also measured the profiles of the beams going into the chamber, somewhere before the 2" folding mirror. They have nice gaussian shapes
 
 
|
300
|
Thu Feb 9 09:30:52 2017 |
Gabriele | Optics | Daily Progress | Two laser installed, setup aligned |
The lasers are behaving well, there is no high noise or wandering lines. The spectrum below is taken in air: that explains the excess of noise in the few kHz region.

Quote: |
Today I swapped out the 8mW laser and installed two 2.5 mW lasers. I rebuilt the input part of the optical levers and re-aligned everything. See below for a picture of the new setup: red beams are input, yellow beams are output. I also installed a protective screen all around the table, to abvoid any suprios beam to get out.

|
|
299
|
Thu Feb 9 09:10:17 2017 |
Gabriele, Larry | Electronics | Configuration | Private network |
I reconfigured the powerstrip and the two Newport controllers to use static IP address in the new 10.10.10.X network, using the workstation as gateway. All scripts updated and working
Quote: |
This morning Larry set up a network switch to create a local network for the power strip and the two Newport picomotor controllers. The laboratory workstation serves as gateway between the networks.
I still have to configure the power strip and the picomotor controllers to use the new static IP address.
|
|
298
|
Thu Feb 9 08:35:24 2017 |
Gabriele | Electronics | Configuration | Removed ESD bias path to model CR0 |
Removed, since iit was not useful to improve the actuation authority.
Quote: |
I added to the model CR0 an additional bias path for the ESD driver:
|
|
297
|
Wed Feb 8 17:06:11 2017 |
Gabriele, Larry | Electronics | Configuration | Private network |
This morning Larry set up a network switch to create a local network for the power strip and the two Newport picomotor controllers. The laboratory workstation serves as gateway between the networks.
I still have to configure the power strip and the picomotor controllers to use the new static IP address. |
296
|
Wed Feb 8 17:01:01 2017 |
Gabriele | Optics | Daily Progress | Two laser installed, setup aligned |
Today I swapped out the 8mW laser and installed two 2.5 mW lasers. I rebuilt the input part of the optical levers and re-aligned everything. See below for a picture of the new setup: red beams are input, yellow beams are output. I also installed a protective screen all around the table, to abvoid any suprios beam to get out.

|
295
|
Tue Feb 7 16:16:37 2017 |
Gabriele | Optics | Daily Progress | Lasers |
The high power lasers I tested so far (the Thorlabs 21mW and the JDSU 1125P) are noisy: they both have wandering lines that from time to time are alised down into the base band, destroing the measurement.
I have three JDSU 1103P units: two of them dlived about 2.5 mW, the third one delivers about 1.4 mW. One of the 2.5mW was installed in the test setup. I swapped it out with the 1.4 mW, so now I have two good 2.5 mW laser. My plan is to modify the new setup to use those two lasers in parallel, splitting each one in two, for a total of four beams of about 1.2 mW each.
The new optical layout is atttached. |
294
|
Tue Feb 7 10:03:35 2017 |
Gabriele | General | Measurements | S1600496 |
2017-02-07
- 10:02am in chamber, balanced
- 10:03am roughing pump on
- 10:30am turbo pump on
- Excitations (they didn't work, I forgot to switch on the noise!!)
-
Quiet time before excitation: 1170548335
Excitation (broad band) at 1170548367 (60 s)
Quiet time after excitation: 1170548429
-
Quiet time before excitation: 1170559259
Excitation (broad band) at 1170559291 (60 s)
Quiet time after excitation: 1170559353
-
Quiet time before excitation: 1170570183
Excitation (broad band) at 1170570216 (60 s)
Quiet time after excitation: 1170570278
-
Quiet time before excitation: 1170581108
Excitation (broad band) at 1170581140 (60 s)
Quiet time after excitation: 1170581202
-
Quiet time before excitation: 1170592032
Excitation (broad band) at 1170592064 (60 s)
Quiet time after excitation: 1170592126
-
Quiet time before excitation: 1170602956
Excitation (broad band) at 1170602989 (60 s)
Quiet time after excitation: 1170603051
-
Excitations:
-
Quiet time before excitation: 1170638328
Excitation (broad band) at 1170638360 (60 s)
Quiet time after excitation: 1170638422
-
Quiet time before excitation: 1170649252
Excitation (broad band) at 1170649285 (60 s)
Quiet time after excitation: 1170649347
-
Quiet time before excitation: 1170660177
Excitation (broad band) at 1170660209 (60 s)
Quiet time after excitation: 1170660271
-
Quiet time before excitation: 1170671101
Excitation (broad band) at 1170671133 (60 s)
Quiet time after excitation: 1170671195
-
Quiet time before excitation: 1170682025
Excitation (broad band) at 1170682057 (60 s)
Quiet time after excitation: 1170682119
2017-02-09
|
293
|
Mon Feb 6 16:22:47 2017 |
Gabriele | General | Measurements | S1600472 S1600478 S1600473 S1600480 |
- In the new chamber
- Excitations:
-
Quiet time before excitation: 1170443490
Excitation (broad band) at 1170443523 (60 s)
Quiet time after excitation: 1170443586
-
Quiet time before excitation: 1170454416
Excitation (broad band) at 1170454449 (60 s)
Quiet time after excitation: 1170454512
-
Quiet time before excitation: 1170465342
Excitation (broad band) at 1170465375 (60 s)
Quiet time after excitation: 1170465438
2017-02-07
- 9:36am gate valve closed, pumps off
|
292
|
Mon Feb 6 14:09:07 2017 |
Gabriele | General | Measurements | First set of parallel measurements completed |
The new laser is still acting up a bit, with a wandering line moving sometimes inside the band. Nevertheless, I could perform the first complete measurement of four samples in parallel.
Everything went as expected, there is no visible cross coupling between modes of different disks. The MATLAB code for the analysis is ready and working.
Here are the results:
 
 
|
291
|
Mon Feb 6 11:19:13 2017 |
Gabriele | Optics | Daily Progress | Laser swap |
The Thorlabs laser has been misbehaving for the whole weekend. Even after many days being continuosly on, the wandering line is still moving all over the frequencies.
So this morning I swapped in a JDSU 1125P borrowed from the 40m lab, which provides about 6.8 mW of power. I tested it over the weekend on a separate test table, and after one day or so of operation the power looks reasonably stable. Now it's been on for a few hours: there is still a line moving around, but it's slowing down and hopefully setting down in a good place.
I started a series of test measurements on the samples that were already installed.
- Quiet time before excitation: 1170443490
Excitation (broad band) at 1170443523 (60 s)
Quiet time after excitation: 1170443586
|
290
|
Mon Feb 6 08:37:54 2017 |
Gabriele | General | Measurements | S1600495 |
2017-02-06
- 8:35am in chamber, balanced
- 8:36am roughing pump on
- 8:50am tubo pump on
- Excitations: (they didn't work, I forgot to switch on the noise!!)
-
Quiet time before excitation: 1170460053
Excitation (broad band) at 1170460085 (60 s)
Quiet time after excitation: 1170460147
-
Quiet time before excitation: 1170470977
Excitation (broad band) at 1170471010 (60 s)
Quiet time after excitation: 1170471072
-
Quiet time before excitation: 1170481902
Excitation (broad band) at 1170481935 (60 s)
Quiet time after excitation: 1170481997
-
Quiet time before excitation: 1170492827
Excitation (broad band) at 1170492860 (60 s)
Quiet time after excitation: 1170492922
-
Quiet time before excitation: 1170503752
Excitation (broad band) at 1170503785 (60 s)
Quiet time after excitation: 1170503847
-
Quiet time before excitation: 1170514677
Excitation (broad band) at 1170514710 (60 s)
Quiet time after excitation: 1170514772
2017-02-07
|
289
|
Wed Feb 1 16:58:10 2017 |
Gabriele | General | Characterization | Excitation tests |
The excitation is working fine. However, there is a large cross coupling in the excitation between channels: so if I switch on only one of the ESD paths, I actually excite all other disks too. It might be due to the common ground or to the fact that the cables run close to each other. This needs some investigation, but it's not a big issue for the moment being.
On the other hand, the optical levers are very well decoupled: each one sees a different set of modes. So there is no measurable cross coupling between the disks or the readout.
There seems to be something fishy with the Y picomotor of QPD3: it doesn't always move in the same direction for the same set of steps. Some investigations needed here too. Autocentering might fail, but it's still possible to center it manually.
Excitation at 4:56pm (quiet time before excitation 1170032165, quiet time after excitation 1170032234).
Another excitation at 8:30am (2017-02-02) (quiet time before excitation 1170088185, quiet time after excitation 1170088373) |
288
|
Wed Feb 1 15:35:30 2017 |
Gabriele | General | Daily Progress | excitation not working - FIXED |
My bad, a cable was disconnected.
Quote: |
Finally, I tried to excite the disks, but I couldn't get any motion of the optical lever beams. To be investigated, there might simply be some cables disconnected.
|
|
287
|
Wed Feb 1 15:23:42 2017 |
Gabriele | Optics | Noise hunting | Wandering line is due to the laser |
The wandering line I mentioned in my previous elog, which is spoiling most of the sensitivity, turns out to be power noise of the laser.
I used a Thorlabs PDA100 and a SR785 to measure the power noise out of the laser directly, and saw a huge forest of peaks above 20kHz. Among them, a couple of peaks are moving up and down in frequency very fast. The plot below compares two different times of the Thorlabs HNL210L laser (the new one, 21 mW) with the old JTSU laser we are using for the test setup:

The noise of the new laser is cleary much larger (even after the laser has been on for some time) and non stationary. This is a big issue for us. I will contact Thorlabs to inquire if this behavior is normal.
The attached video file shows the peaks dancing around on the SR785 screen. |
286
|
Tue Jan 31 16:55:17 2017 |
Gabriele | General | Daily Progress | Improved autocenter, excitation not working, wandering line |
The autocenter script wasn't working in a very robust way (sometimes the socket connection to the controller failed, some times no motion was obtained after issuing a command). So I rewrote the interface to the Newport controller using the http server interface. The code is in picomotor8742_http.py. This version seems much more robust.
There is a large wandering line that spans all frequencies. Not sure what the origin is, but it will need some noise hunting. Here are spectrograms of all four QPD X signals. The line is visible in all of them, at the same frequency. So it's definitely something external coupling into the light or the QPD electronics. It's not visible in the sum of the QPD quadrants, but maybe just because it's buried in noise.

Finally, I tried to excite the disks, but I couldn't get any motion of the optical lever beams. To be investigated, there might simply be some cables disconnected.
|
285
|
Tue Jan 31 16:50:55 2017 |
Gabriele, Alena | General | Vacuum | First pump-down with setup |
This afternoon we started the pump-down with all the system installed into the chamber. Unfortunately the IGM vacuum gauge isn't working, so we can't be sure what the pressure is. To be fixed |
284
|
Tue Jan 31 11:21:44 2017 |
Gabriele | Mechanics | Daily Progress | Retaining ring mechanics |
All parts for the motion of the retaining rings have been received and are ok. We're going to clean and bake them.
 
|
283
|
Tue Jan 31 09:03:12 2017 |
Gabriele | Electronics | Configuration | Auto centering script for the new setup |
I wrote an autocenter script for the four QPD in the new setup (autocenter14.py) and renamed the script for the old chamber (autocenter0.py).
Tested and working properly, with network connection to the picomotor controllers. Be aware that if the picomotor controllers are switched off, their IP address might change. |
282
|
Mon Jan 30 15:16:54 2017 |
Gabriele | General | Daily Progress | Four samples in chamber |
- checked that the high voltage paths are working:
- the DAC signal is properly received and amplified by the Trek 2220 HV amplifier
- the HV switch box is working as expected: I can switch on and off each path individiually using the DAC signals
- installed four samples (refer to the optical drawing for the numbering)
- S1600472 in bay 1
- S1600478 in bay 2
- S1600473 in bay 3
- S1600480 in bay 4
 
- balanced all samples and centered all optical levers. There is a problem with the fifth connector in the ADC interface board (or maybe with one ADC channel). So one of the quadrants of QPD4 was not working. I moved the connector to number 6 and updated the model. Everything looks fine now.
- the power in the four QPD is quite diverse: this is not unexpected, since it depends on the interference of the reflection from the two surfaces of the disk

- installed the picomotor controllers on the bench, connected them to the net using the switch on the top of the clean room. For the moment being they are using DHCP.
- connected all picomotors to the controllers.
- Bottom controller, from left to right: QPD1x, QPD1y, QPD2x, QPD2y
- Top controller, from left to right: QPD3x, QPD3y, QPD4x, QPD4y
- tested the remote control and motion of all picomotors
  |